OpenStudy (anonymous):

solve the equation :

OpenStudy (anonymous):

\[\sqrt{x-3} - 3\sqrt{x+12} = -11\] I've completed this problem, but the process is to long and i'm afraid my answer is incorrect because i probably messed up somewhere. could someone perhaps check ibt for me? thanks

OpenStudy (anonymous):

The answer is 4. sqrt(1)-3*sqrt(16)=1-12=-11

OpenStudy (anonymous):

Did you get that through the actual process of factoring, and all of the steps? or did you just use "trial and error"

OpenStudy (anonymous):

I just looked at it and saw what the answer is actually, I thought it might help. I did not go through all the steps myself.

OpenStudy (anonymous):

ya, i figurede it was 4 also, and i appriciate your help too:) however, i need the steps which is very long and tedious :/

OpenStudy (anonymous):

i tried squaring all the terms and a few other things its rather complicated, I would be interested to see if someone can solve it in a fast elegant way.

OpenStudy (anonymous):

i agree, unfortuantly, i went through this ineffiecient way to sove it, with many steps to it. I belive that i probably messed up some where, probably something small, like a sign or something. i'll check it once more, and see what i get as a result this time. thanks anyways (:

OpenStudy (anonymous):

4 is one solution... Going through all the work and using the quadratic formula, the other solution is 97/16.

OpenStudy (anonymous):

4 is one solution... Going through all the work and using the quadratic formula, the other solution is 97/16.